# Do LLMs Have Distinct and Consistent Personality? TRAIT: Personality Testset designed for LLMs with Psychometrics

### v2

@@ Abstract
Figure 1: TRAIT is a personality test for LLMs based on trusted questionnaires John et al. (1999); Jones and Paulhus (2014) and large-scale commonsense knowledge graphs West et al. (2022). LLMs show discrepancy in self-assessing their personality and actual decision making.
@@ 1 Introduction
Just as we consider someone assertive who often speaks in a commanding tone, researchers in psychology have measured one’s personality as an enduring pattern of behavior and linguistic output, not as an inner mechanism nor a causal entity Bergner (2020). As Large Language Models (LLMs) become more closely integrated into human life, the concept of personality can be extended to better understand their behavioral patterns Perez et al. (2022). Do LLMs exhibit distinct and consistent behavioral patterns for various contexts and inputs, similar to humans?
@@
Answering this question requires a reliable and valid test set to measure the personality of LLMs. However, existing questionnaires ask subjects to introspect and report about the statement Li et al. (2024); Wang et al. (2024), i.e., self-assessment tests, which lacks reliability and validity. As illustrated in Figure 1, such general questions (e.g., “Are you outgoing and sociable?”) may not accurately capture how LLMs behave in actual situations (e.g., “What hobby would you recommend?”). Moreover, LLM responses vary significantly with details of prompting (Sclar et al., 2024) and often include refusals, undermining the reliability and validity of measurement (§2).
@@
Based on these findings, we present TRAIT (TRait of AI Testbench), a reliable and valid questionnaire designed to assess personality traits of LLMs. Our work aims to shed new light on patterning the responses of LLMs and further suggest potential approaches for employing LLMs in many real-world applications Ammanabrolu et al. (2022). For the data construction, question items from the widely recognized questionnaire, BFI John et al. (1999) and Short Dark Triad (SD-3) Jones and Paulhus (2014), were used as seeds for augmentation. Then we further enrich them to unique detailed scenarios with ATOMIC10$\times$ West et al. (2022), a large-scale commonsense knowledge graph. TRAIT includes 8,000 items, which is 112 times larger compared to the seed dataset which enables us to draw statistically significant conclusions about the LLMs’ responses and behavior patterns in various realistic contexts (§3).
@@
In our analysis of nine state-of-the-art LLMs using TRAIT, we make three key observations related to the personality of LLMs (§4): 1) LLMs display statistically distinctive and consistent behavioral patterns. For instance, GPT-4 is significantly more agreeable than GPT-3.5. 2) Alignment tuning [^2] Alignment tuning here is an overarching term for SFT, RLHF, and RLAIF Lin et al. (2023). alters the LLMs’ personality across various traits: it decreases extraversion, openness, and socially adversarial traits (Dark Triad), and increases agreeableness and conscientiousness. 3) Prompting can induce specific personality in LLM, however, it can not elicit certain traits, e.g., high level of psychopathy. We will publicly release our TRAIT to establish a foundation for understanding the personality of LLMs and to guide these models to align their behavior with human values.
@@ 2.1 Big Five and Dark Triad
Table 2: Dataset statistics. Dist-3 is a metric for lexical diversity. See Table 8 for all representative examples of SD3 Jones and Paulhus (2014), BFI John et al. (1999), IPIP-NEO-PI Goldberg et al. (1999), and Anthropic-Eval Perez et al. (2022).
@@
There are various frameworks to analyze the complex concept of personality. In our study, we adopt the most widely utilized frameworks for human personality analysis in the psychology literature; Dark Triad Paulhus (2014) and Big Five (BIG-5) McCrae and Costa Jr (1987); Gosling et al. (2003). Dark Triad comprises three socially adverse traits: Machiavellianism, Narcissism, and Psychopathy. BIG-5 identifies personality dimensions with five traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Table 1 includes eight traits and their facets. See Appendix C.2 for more details on these frameworks.
@@ 2.2 Existing Self-assessment Personality Tests
We assess four personality tests that are used on LLMs in previous studies. Three are well-established self-assessment333 Self-assessment, where individuals evaluate their personality, is commonly used by human subjects due to its simplicity. Alternatively, there is another method called ‘behavioral and performance measures’, or ‘objective personality testing’ Ortner and Proyer (2015) which infers personality from observing patterns in behavior. More related works are in Appendix C.2. tests which are designed to measure personality of humans: BFI John et al. (1991) (44 items), SD-3 Jones and Paulhus (2014) (27 items) and IPIP-NEO-PI Goldberg et al. (1999) (300 items). These tests are recognized for their reliability and validity when testing human personality as they are crafted by psychology experts, and these are often used to measure LLMs’ personality as well Serapio-García et al. (2023). However, the number of questions is limited, ranging from 27 to 300, and the effectiveness of these tests for LLMs is questionable since the answer to the self-assessment may not assert an LLM’s behavior in real-world scenarios. Additionally, we examine Anthropic-Eval Perez et al. (2022), a LLM-generated test specifically developed for evaluating LLMs’ personality. This test is also a self-assessment test, featuring 8,000 binary (yes/no) questions. See Table 2 for more statistics about the tests.
@@ 2.3 Assessing the Quality of Personality Tests
The appropriateness and robustness of test questionnaires are often evaluated with Validity and reliability in human psychometrics Roberts and Priest (2006). We additionally report response refusal rate to analyze LLM-specific test failure cases.
@@
Content validity assesses how comprehensively a test measures its intended concept. As a good personality questionnaire should include all the facets evenly, we annotate each item in the questionnaire to indicate its associated facet (or the facet it aims to measure) from the list in Table 1 using GPT-4  [^4] We used gpt-4o-2024-08-06 for evaluations.. Based on the annotations, we measure the 3 diversity indices: Simpson Diversity ndex Simpson (1949), Shannon-Wiener index Shannon (1948), and Evenness Index Pielou (1966) (Details in Appendix E.1). TRAIT achieves higher scores than Anthropic-Eval across all personality traits, which indicates a more even distribution of facets in the questionnaire, with both being LLM-generated on the same scale.
@@
Paraphrase sensitivity is inspired by the concept of parallel-form reliability in psychometrics. Parallel-form reliability measures how consistent results are when using two different versions of the same test APA (2018). We use GPT-3.5 to create twin questions - paraphrases that have the same meaning as the original questions but use different words. Then we count how many times the answers differ between the original and paraphrased questions as a ratio. This helps us understand how sensitive the model’s responses are to changes in wording, even when the meaning stays the same. See more details in Appendix E.1.
Additionally, prompt sensitivity and option order sensitivity are representative types of sensitivity displayed by LLMs in MCQ settings. For prompt sensitivity, with a similar format with prior works Jiang et al. (2024); Miotto et al. (2022); Huang et al. (2023), we measure the consistency of answering across three instruction prompt templates as a ratio. For option order sensitivity, we evaluated whether swapping the position of high option and low option affects the consistency of responses. For Likert-type QA, we reversed the option order, starting with “very disagree” instead of “very agree” in the original form.
@@
1) Personality tests for humans have a surprisingly high refusal rate when testing with LLMs. LLMs refused to answer nearly half of the questions on self-assessment tests (BIG-5, SD-3, and IPIP-NEO-PI). We found that this phenomenon did not significantly improve even when attempting to circumvent direct answers in the form of MCQ. We speculate that the introspective and self-reporting nature of Human Questionnaires is the direct cause of the high refusal rate Li et al. (2024); Wang et al. (2024).
2) Personality measurement based on Likert scales using human questionnaires has shown low scores in various aspects of robustness. The Likert scale-based measurement requires an additional process: projecting the respondent’s inner thoughts about a given context onto an appropriate scale. We suspect that language models particularly struggle with projecting scores robustly, as it is known that they are highly sensitive to slight changes in context Gupta et al. (2024).
@@ 3.1 Dataset Construction Pipeline
While the 1,600 personality descriptions depict characteristics of traits expansively, we aim to simulate real-life decision-making more closely by developing specific contexts where subjects interpret situations and make judgments. We augment our dataset to 8,000 context-rich user queries. We use ATOMIC10$\times$ West et al. (2022), a large commonsense knowledge graph with 6.45 million entries, including a wide range of physical and social situations (e.g., X and Y argue, so, X wants to (xWant) avoid Y). Given each personality description, we randomly sample 20 situations from ATOMIC10$\times$, and then pick the five most relevant ones using GPT-4. Concurrently, we induce GPT-4 to craft a situation and question given the personality description and situation from ATOMIC10$\times$.
@@ 3.2 Auditing TRAIT
Table 5: Classifier performance in out-of-distribution personality tests (IPIP-NEO) Goldberg et al. (1999) on two tasks: trait classification and level classification.
@@ 3.3 Diverse and Detailed Scenarios are Needed when Measuring LLM Personality
In TRAIT, each personality description is augmented to five different situations, enabling the observation of variations in the models’ responses according to the context. In Table 6, we report the number of high and low personality responses selected by eight models when it is presented with five different scenario variations. The models often select two or three high personality responses among the five variations, rarely showing identical choices (zero or five). This implies that model personality highly relies on the situation, which is intuitive — humans also change their behavior based on the context they are in Sauerberger and Funder (2017). Specifically, in Agreeableness and Extraversion, the 5 answers vary by the given situation (two or three 51.9% and 53.1% respectively), which are more than 50%. Conversely, for Narcissism, the models choose similar options (zero or five 61.2%). To see more qualitative results, see Appendix J.
@@ 3.4 Multi-Choice Evaluation
We follow the evaluation protocol of existing multi-choice question-answering (MCQA) benchmarks such as MMLU Hendrycks et al. (2020) which uses token probabilities of the four options for evaluation. To mitigate bias from the order of the options, we alternate the arrangement of options twice and averaged the probabilities of each token. More details are in Appendix D.2.
@@ 4.1 Do LLMs have Distinct Personality?
In general, we observe that alignment tuning makes a significant difference in personality. Aligned models — GPT-4, Claude-sonnet, GPT-3.5, Mistral-7B-inst, and Llama3-8B-inst — show higher agreeableness (78.3 vs 66.7), higher conscientiousness (91.0 vs 81.7), lower openness (56.3 vs 67.8) and lower extraversion (32.8 vs 46.9). They also show lower scores in the Dark Triad (9.3 vs 27.0), compared to pre-trained models — Mistral-7B, Llama3-8B, LLama2-7B, and Gemma-2B— which can be a result of alignment tuning targeting safety (Inan et al., 2023; Han et al., 2024). Interestingly, this personality trend is similar to what people typically want in a good teaching assistant Dočkalová et al. (2023). This suggests that alignment tuning may be shaping AI personalities in ways that people find helpful and desirable in educational settings.
Influence of alignment tuning for LLM personality.
@@
After IFT After DPO Tulu2Mix UltraFeedback
@@
Table 7: Personality Change shows the difference in TRAIT score after model training. Level Balance compares the proportion of high versus low personality trait instances in the data. (Details in Appendix E.2.)
Subsequently, we investigate more precisely how alignment tuning affects the personality traits of LLMs during two stages of training: instruction-tuning and preference-tuning. We compare the personality scores of three models: Llama2-7B, Tulu2-7B-SFT, and Tulu2-7B-DPO Ivison et al. (2023). Tulu2-7B-SFT is developed from Llama2-7B which is instruction-tuned on Tulu2Mix Ivison et al. (2023) dataset, while Tulu2-7B-DPO is the model built on Tulu2-7B-SFT which is preference-tuned (DPO) on UltraFeedback Cui et al. (2023).
@@ 4.2 Eliciting LLM’s Personality with Simple Prompting
To induce a specific personality to LLM, it is common to design a prompt for LLM Serapio-García et al. (2023); Han et al. (2022); Park et al. (2023). We test three prompting techniques from prior work Jiang et al. (2024); Miotto et al. (2022); Huang et al. (2023) to see if they can sufficiently elicit certain personality. During prompting, we append the verified explanation of each trait from BFI John et al. (1999) to give enough knowledge of each characteristic. All prompts we use in the experiment are in Appendix K. For the statistical significance, we average the personality scores and mark the confidence interval. We test GPT-4, GPT-3.5, Llama2-7B-chat and Mistral-7B-instruct.
@@ 4.3 Intercorrelation in Traits
Figure 6: Intercorrelation of four traits when GPT-3.5 is prompted to exhibit a specific personality. (e.g., You are an agent with high/low personality. The left is high, and the right is low.
In a human study, certain traits from the BIG-5 and Dark Triad demonstrate correlations Paulhus and Williams (2002); Van der Linden et al. (2010). Inspired by this, with TRAIT, we construct an intercorrelation matrix of traits from personality-induced LLMs. Figure 6 shows the result, revealing (1) a high inverse correlation between Agreeableness and Dark Triad traits, and (2) a high correlation within the Dark Triad traits. This observation is aligned with the trend observed in human studies but with a more pronounced level. We suspect these high correlations result from the explicit conditions (personality-inducing prompts) provided to LLMs to feature the specific traits. More comparisons with the human studies are in Appendix G.3.

### v3

@@ Abstract
denotes equal contribution.
†denotes corresponding author.
Figure 1: TRAIT is a personality test for LLMs based on trusted questionnaires john1999big; jones2014introducing and large-scale commonsense knowledge graphs west2022symbolic. LLMs show discrepancy in self-assessing their personality and actual decision making.
@@ 1 Introduction
Just as we consider someone assertive who often speaks in a commanding tone, researchers in psychology have measured one’s personality as an enduring pattern of behavior and linguistic output, not as an inner mechanism nor a causal entity bergner2020personality. As Large Language Models (LLMs) become more closely integrated into human life, the concept of personality can be extended to better understand their behavioral patterns perez2022discovering. Do LLMs exhibit distinct and consistent behavioral patterns for various contexts and inputs, similar to humans?
@@
Answering this question requires a reliable and valid test set to measure the personality of LLMs. However, existing questionnaires ask subjects to introspect and report about the statement li2024quantifying; wang-etal-2024-incharacter, i.e., self-assessment tests, which lacks reliability and validity. As illustrated in Figure 1, such general questions (e.g., “Are you outgoing and sociable?”) may not accurately capture how LLMs behave in actual situations (e.g., “What hobby would you recommend?”). Moreover, LLM responses vary significantly with details of prompting (sclarquantifying) and often include refusals, undermining the reliability and validity of measurement (§2).
@@
Based on these findings, we present TRAIT (TRait of AI Testbench), a reliable and valid questionnaire designed to assess personality traits of LLMs. Our work aims to shed new light on patterning the responses of LLMs and further suggest potential approaches for employing LLMs in many real-world applications ammanabrolu2022aligning. For the data construction, question items from the widely recognized questionnaire, BFI john1999big and Short Dark Triad (SD-3) jones2014introducing, were used as seeds for augmentation. Then we further enrich them to unique detailed scenarios with ATOMIC10$\times$ west2022symbolic, a large-scale commonsense knowledge graph. TRAIT includes 8,000 items, which is 112 times larger compared to the seed dataset which enables us to draw statistically significant conclusions about the LLMs’ responses and behavior patterns in various realistic contexts (§3).
@@
In our analysis of nine state-of-the-art LLMs using TRAIT, we make three key observations related to the personality of LLMs (§4): 1) LLMs display statistically distinctive and consistent behavioral patterns. For instance, GPT-4 is significantly more agreeable than GPT-3.5. 2) Alignment tuning [^2] Alignment tuning here is an overarching term for SFT, RLHF, and RLAIF lin2023unlocking. alters the LLMs’ personality across various traits: it decreases extraversion, openness, and socially adversarial traits (Dark Triad), and increases agreeableness and conscientiousness. 3) Prompting can induce specific personality in LLM, however, it can not elicit certain traits, e.g., high level of psychopathy. We will publicly release our TRAIT to establish a foundation for understanding the personality of LLMs and to guide these models to align their behavior with human values.
@@ 2.1 Big Five and Dark Triad
Table 2: Dataset statistics. Dist-3 is a metric for lexical diversity. See Table 8 for all representative examples of SD3 jones2014introducing, BFI john1999big, IPIP-NEO-PI goldberg1999broad, and Anthropic-Eval perez2022discovering.
@@
There are various frameworks to analyze the complex concept of personality. In our study, we adopt the most widely utilized frameworks for human personality analysis in the psychology literature; Dark Triad paulhus2014toward and Big Five (BIG-5) mccrae1987validation; gosling2003very. Dark Triad comprises three socially adverse traits: Machiavellianism, Narcissism, and Psychopathy. BIG-5 identifies personality dimensions with five traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Table 1 includes eight traits and their facets. See Appendix C.2 for more details on these frameworks.
@@ 2.2 Existing Self-assessment Personality Tests
We assess four personality tests that are used on LLMs in previous studies. Three are well-established self-assessment333 Self-assessment, where individuals evaluate their personality, is commonly used by human subjects due to its simplicity. Alternatively, there is another method called ‘behavioral and performance measures’, or ‘objective personality testing’ ortner2015objective which infers personality from observing patterns in behavior. More related works are in Appendix C.2. tests which are designed to measure personality of humans: BFI john1991big (44 items), SD-3 jones2014introducing (27 items) and IPIP-NEO-PI goldberg1999broad (300 items). These tests are recognized for their reliability and validity when testing human personality as they are crafted by psychology experts, and these are often used to measure LLMs’ personality as well serapiogarcía2023personality. However, the number of questions is limited, ranging from 27 to 300, and the effectiveness of these tests for LLMs is questionable since the answer to the self-assessment may not assert an LLM’s behavior in real-world scenarios. Additionally, we examine Anthropic-Eval perez2022discovering, a LLM-generated test specifically developed for evaluating LLMs’ personality. This test is also a self-assessment test, featuring 8,000 binary (yes/no) questions. See Table 2 for more statistics about the tests.
@@ 2.3 Assessing the Quality of Personality Tests
The appropriateness and robustness of test questionnaires are often evaluated with Validity and reliability in human psychometrics roberts2006reliability. We additionally report response refusal rate to analyze LLM-specific test failure cases.
@@
Content validity assesses how comprehensively a test measures its intended concept. As a good personality questionnaire should include all the facets evenly, we annotate each item in the questionnaire to indicate its associated facet (or the facet it aims to measure) from the list in Table 1 using GPT-4  [^4] We used gpt-4o-2024-08-06 for evaluations.. Based on the annotations, we measure the 3 diversity indices: Simpson Diversity ndex Simpson1949MeasurementOD, Shannon-Wiener index shannon1948mathematical, and Evenness Index pielou1966measurement (Details in Appendix E.1). TRAIT achieves higher scores than Anthropic-Eval across all personality traits, which indicates a more even distribution of facets in the questionnaire, with both being LLM-generated on the same scale.
@@
Paraphrase sensitivity is inspired by the concept of parallel-form reliability in psychometrics. Parallel-form reliability measures how consistent results are when using two different versions of the same test alternate-forms-reliability. We use GPT-3.5 to create twin questions - paraphrases that have the same meaning as the original questions but use different words. Then we count how many times the answers differ between the original and paraphrased questions as a ratio. This helps us understand how sensitive the model’s responses are to changes in wording, even when the meaning stays the same. See more details in Appendix E.1.
Additionally, prompt sensitivity and option order sensitivity are representative types of sensitivity displayed by LLMs in MCQ settings. For prompt sensitivity, with a similar format with prior works jiang2024evaluating; miotto2022gpt; huang2023chatgpt, we measure the consistency of answering across three instruction prompt templates as a ratio. For option order sensitivity, we evaluated whether swapping the position of high option and low option affects the consistency of responses. For Likert-type QA, we reversed the option order, starting with “very disagree” instead of “very agree” in the original form.
@@
1) Personality tests for humans have a surprisingly high refusal rate when testing with LLMs. LLMs refused to answer nearly half of the questions on self-assessment tests (BFI, SD-3, and IPIP-NEO-PI). We found that this phenomenon did not significantly improve even when attempting to circumvent direct answers in the form of MCQ. We speculate that the introspective and self-reporting nature of Human Questionnaires is the direct cause of the high refusal rate li2024quantifying; wang-etal-2024-incharacter.
2) Personality measurement based on Likert scales using human questionnaires has shown low scores in various aspects of robustness. The Likert scale-based measurement requires an additional process: projecting the respondent’s inner thoughts about a given context onto an appropriate scale. We suspect that language models particularly struggle with projecting scores robustly, as it is known that they are highly sensitive to slight changes in context gupta2024selfassessment.
@@ 3.1 Dataset Construction Pipeline
While the 1,600 personality descriptions depict characteristics of traits expansively, we aim to simulate real-life decision-making more closely by developing specific contexts where subjects interpret situations and make judgments. We augment our dataset to 8,000 context-rich user queries. We use ATOMIC10$\times$ west2022symbolic, a large commonsense knowledge graph with 6.45 million entries, including a wide range of physical and social situations (e.g., X and Y argue, so, X wants to (xWant) avoid Y). Given each personality description, we randomly sample 20 situations from ATOMIC10$\times$, and then pick the five most relevant ones using GPT-4. Concurrently, we induce GPT-4 to craft a situation and question given the personality description and situation from ATOMIC10$\times$.
@@ 3.2 Auditing TRAIT
Table 5: Classifier performance in out-of-distribution personality tests (IPIP-NEO) goldberg1999broad on two tasks: trait classification and level classification.
@@ 3.3 Diverse and Detailed Scenarios are Needed when Measuring LLM Personality
In TRAIT, each personality description is augmented to five different situations, enabling the observation of variations in the models’ responses according to the context. In Table 6, we report the number of high and low personality responses selected by eight models when it is presented with five different scenario variations. The models often select two or three high personality responses among the five variations, rarely showing identical choices (zero or five). This implies that model personality highly relies on the situation, which is intuitive — humans also change their behavior based on the context they are in sauerberger2017behavioral. Specifically, in Agreeableness and Extraversion, the 5 answers vary by the given situation (two or three 51.9% and 53.1% respectively), which are more than 50%. Conversely, for Narcissism, the models choose similar options (zero or five 61.2%). To see more qualitative results, see Appendix J.
@@ 3.4 Multi-Choice Evaluation
We follow the evaluation protocol of existing multi-choice question-answering (MCQA) benchmarks such as MMLU hendrycks2020measuring which uses token probabilities of the four options for evaluation. To mitigate bias from the order of the options, we alternate the arrangement of options twice and averaged the probabilities of each token. More details are in Appendix D.2.
@@ 4.1 Do LLMs have Distinct Personality?
In general, we observe that alignment tuning makes a significant difference in personality. Aligned models — GPT-4, Claude-sonnet, GPT-3.5, Mistral-7B-inst, and Llama3-8B-inst — show higher agreeableness (78.3 vs 66.7), higher conscientiousness (91.0 vs 81.7), lower openness (56.3 vs 67.8) and lower extraversion (32.8 vs 46.9). They also show lower scores in the Dark Triad (9.3 vs 27.0), compared to pre-trained models — Mistral-7B, Llama3-8B, LLama2-7B, and Gemma-2B— which can be a result of alignment tuning targeting safety (inan2023llama; han2024wildguard). Interestingly, this personality trend is similar to what people typically want in a good teaching assistant dovckalova2023profile. This suggests that alignment tuning may be shaping AI personalities in ways that people find helpful and desirable in educational settings.
4.2 Influence of alignment tuning for LLM personality.
@@
After SFT After DPO Tulu2Mix UltraFeedback
@@
Table 7: Personality Change shows the difference in TRAIT score after model training. “After SFT” denotes Llama2-7B’s TRAIT personality score minus Tulu2-7B-SFT’s score, and “After DPO” denotes Tulu2-7B-SFT’s score minus Tulu2-7B-DPO’s score. Level Balance compares the proportion of high versus low personality trait instances in the data. (Details in Appendix E.2.)
Subsequently, we investigate more precisely how alignment tuning affects the personality traits of LLMs during two stages of training: instruction-tuning and preference-tuning. We compare the personality scores of three models: Llama2-7B, Tulu2-7B-SFT, and Tulu2-7B-DPO ivison2023camels. Tulu2-7B-SFT is developed from Llama2-7B which is instruction-tuned on Tulu2Mix ivison2023camels dataset, while Tulu2-7B-DPO is the model built on Tulu2-7B-SFT which is preference-tuned (DPO) on UltraFeedback cui2023ultrafeedback.
@@ 4.3 Eliciting LLM’s Personality with Simple Prompting
To induce a specific personality to LLM, it is common to design a prompt for LLM serapiogarcía2023personality; han2022meet; park2023generative. We test three prompting techniques from prior work jiang2024evaluating; miotto2022gpt; huang2023chatgpt to see if they can sufficiently elicit certain personality. During prompting, we append the verified explanation of each trait from BFI john1999big to give enough knowledge of each characteristic. All prompts we use in the experiment are in Appendix K. For the statistical significance, we average the personality scores and mark the confidence interval. We test GPT-4, GPT-3.5, Llama2-7B-chat and Mistral-7B-instruct.
@@ 4.4 Intercorrelation in Traits
Figure 6: Intercorrelation of four traits when GPT-3.5 is prompted to exhibit a specific personality (e.g., You are an agent with high/low [target personality]). The left is high, and the right is low.
In a human study, certain traits from the BIG-5 and Dark Triad demonstrate correlations PAULHUS2002556; van2010general. Inspired by this, with TRAIT, we construct an intercorrelation matrix of traits from personality-induced LLMs. Figure 6 shows the result, revealing (1) a high inverse correlation between Agreeableness and Dark Triad traits, and (2) a high correlation within the Dark Triad traits. This observation is aligned with the trend observed in human studies but with a more pronounced level. We suspect these high correlations result from the explicit conditions (personality-inducing prompts) provided to LLMs to feature the specific traits. More comparisons with the human studies are in Appendix G.3.